Find three consecutive numbers whose sum is 60
Answer by ptaylor(1894)

Let x=first number
Then (x+1)=second number
And (x+2)=third number
And we are told that:
x+(x+1)+(x+2)=60 get rid of parens
x+x+1+x+2=60 collect like terms
3x+3=60 subtract 3 from both sides
3x+3-3=60-3 simplify
3x=57 divide both sides by 3
x=19--------------------------------first number
x+1=19+1=20-----------------------------second number
x+2=19+2=21------------------------------third number
CK
19+20+21=60
60=60

Hope this helps-----ptaylor